OpenStudy (anonymous):

Andrea Gilford's savings account has a balance of $ 169. After 2 years, the amount of interest at 12% compounded quarterly will be $

OpenStudy (anonymous):

\[169\times (1+\frac{.12}{4})^{2\times 4}\] or \[169(1.03)^8\] then a calculator